### v2

# A Comprehensive Survey of Dataset Distillation

Abstract
Deep learning technology has developed unprecedentedly in the last decade and has become the primary choice in many application domains. This progress is mainly attributed to a systematic collaboration in which rapidly growing computing resources encourage advanced algorithms to deal with massive data. However, it has gradually become challenging to handle the unlimited growth of data with limited computing power. To this end, diverse approaches are proposed to improve data processing efficiency. Dataset distillation, a dataset reduction method, addresses this problem by synthesizing a small typical dataset from substantial data and has attracted much attention from the deep learning community. Existing dataset distillation methods can be taxonomized into meta-learning and data matching frameworks according to whether they explicitly mimic the performance of target data. Although dataset distillation has shown surprising performance in compressing datasets, there are still several limitations such as distilling high-resolution data. This paper provides a holistic understanding of dataset distillation from multiple aspects, including distillation frameworks and algorithms, factorized dataset distillation, performance comparison, and applications. Finally, we discuss challenges and promising directions to further promote future studies on dataset distillation.
1 Introduction
During the past few decades, deep learning has achieved remarkable success in a wide range of applications, including computer vision [57, 41, 19, 42, 20], neural language processing [55, 112, 58], autonomous vehicles [39, 1, 31], and protein structure prediction [100, 51]. An important reason for this success is due to powerful computing resources, which allow deep neural networks to directly handle giant datasets and bypass complicated manual feature extraction, which causes potential loss of data information [109, 110]. For example, the powerful large language model GPT-3 contains $175$ billion parameters and is trained on $45$ terabytes of text data with thousands of graphics processing units (GPUs) [8]. However, massive data are generated every day [97], which pose a significant threat to training efficiency and data storage, and deep learning might reach a bottleneck due to the mismatch between the volume of data and computing resources [104]. Recently, many methods have been proposed to improve training efficiency in deep learning from several perspectives as follows:
  • Quantization: This approach sacrifices data at the byte level during the training process for acceleration [40, 82, 30, 36].
  • Model pruning: This approach removes trainable parameters that have little influence on the final performance [107, 76, 2, 75].
  • Optimization: This approach designs the training algorithms for fast convergence [125] or less memory cost [4, 52].
  • Dataset reduction: this approach generates few representative data to constitute a new training set. Depending on whether the generated data are natural or synthetic, dataset reduction can be classified into coreset selection [83, 22] and dataset distillation [114].
Figure 1: An illustration of dataset distillation. The models trained on the large original dataset and small synthetic dataset demonstrate comparable performance on the test set.
The above narrative shows that efficient deep learning is an extensive task and is beyond the scope of this paper; readers can refer to [108, 81] for more details. In this survey, we focus on dataset distillation to improve training efficiency by synthesizing training data. To better appreciate dataset distillation, we briefly introduce other cognate methods in dataset reduction. For a support vector machine (SVM), its hyperplane is solely determined by “support vectors”, and removing all other points in the training dataset does not influence on the convergence result [23]. Therefore, the selection of “support vectors” is a favorable method to reduce the training burden for SVMs [84]. When the scenario expands to deep learning, the problem of selecting “support vectors” generalizes to the well-known coreset selection; this is the algorithm selects a few prototype examples from the original training dataset as the coreset, and then the model is solely trained on the small coreset to save training costs while avoiding large performance drops. However, elements in the coreset are unmodified and constrained by the original data, which considerably restricts the coreset’s expressiveness, especially when the coreset budget is limited. Recently, a novel approach, dataset distillation (DD)  [^1] Dataset distillation is also referred to as dataset condensation in some studies., has attracted growing attention from the deep learning community and has been leveraged in many practical applications [79, 70, 16, 17]. Different from coreset selection, dataset distillation removes the restriction of uneditable elements and carefully modifies a small number of examples to preserve more information, as shown in Figure 4 for synthetic examples. By distilling the knowledge of the large original dataset into a small synthetic set, models trained on the distilled dataset can acquire a comparable generalization performance. A general illustration for dataset distillation is presented in Figure 1.
Due to the property of extremely high dimensions in the deep learning regime, the data information is hardly disentangled to specific concepts, and thus distilling numerous high-dimensional data into a few points is not a trivial task. Based on the objectives applied to mimic target data, dataset distillation methods can be grouped into meta-learning frameworks and data matching frameworks, and these techniques in each framework can be further classified in a much more detailed manner. In the meta-learning framework, the distilled data are considered hyperparameters and optimized in a nested loop fashion according to the distilled-data-trained model’s risk w.r.t. the target data [77, 114]. The data matching framework updates distilled data by imitating the influence of target data on model training from parameter or feature space [129, 13, 128]. Figure 3 presents these different categories of DD algorithms in a tree diagram.
Figure 2: The schematic structure of dataset distillation and the relationship between the adjacent sections. The body of this survey mainly contains the fundamentals of dataset distillation, taxonomy of distillation schemes, types of factorized $D$, distillation algorithms, performance comparison, applications, challenges, and future directions. Note that ‘Section’ is abbreviated as ‘Sec.’ in this figure.
[…]
Dataset Distillation
Meta-Learning Framework
Backpropagation Through Time
DD [114]
LD [6]
SLDD [106]
GTN [105]
RTP [26]
Kernel Ridge Regression
KIP [86, 87]
FRePo [132]
RFAD [74]
Data Matching Framework
Gradient Match
DC [129]
DCC [63]
IDC [54]
Trajectory Match
MTT [13]
FTD [29]
TESLA [25]
Haba [71]
Distribution Match
DM [128]
CAFE [113]
IT-GAN [127]
KFS [62]
Figure 3: Tree diagram for different categories of dataset distillation algorithms. The factorized DD methods are marked in blue.
Apart from directly considering the synthetic examples as optimization objectives, in some studies, a proxy model is designed, which consists of latent codes and decoders to generate highly informative examples and to resort to learning the latent codes and decoders. For example, Such et al. [105] employed a network to generate highly informative data from noise and optimized the network with the meta-learning framework. Zhao and Bilen [127] optimized the vectors and put them into the generator of a well-trained generative adversarial network (GAN) to produce the synthetic examples. Moreover, Kim et al. [54], Deng and Russakovsky [26], Liu et al. [71], Lee et al. [62] learned a couple of latent codes and decoders, and then synthetic data were generated according to the different combinations of latent codes and decodes. With this factorization of synthetic data, the compression ratio of DD can be further decreased, and the performance can also be improved due to the intraclass information extracted by latent codes.
In this paper, we present a comprehensive survey on dataset distillation, and the main objectives of this survey are as follows: (1) present a clear and systematic overview of dataset distillation; (2) review the recent progress and state-of-the-art algorithms and discuss various applications in dataset distillation; (3) give a comprehensive performance comparison w.r.t. different dataset distillation algorithms; and (4) provide a detailed discussion on the limitation and promising directions of dataset distillation to benefit future studies. Recently, Larasati et al. [59] published a short review on dataset distillation. However, these authors only provided a general overview of DD and focus on the aspect of applications. Different from Larasati et al. [59], our paper gives a systematic, comprehensive survey on dataset distillation from a wide aspect of distillation frameworks, algorithms, applications, limitations, and promising directions.
The rest of this paper is organized as follows. We first provide soem background with respect to DD in Section 2. DD methods under the meta-learning framework are described and comprehensively analyzed in Section 3, and a description of the data matching framework follows in Section 4. Section 5 discusses different types of factorized dataset distillation. Next, we report the performance of various distillation algorithms in Section 6. Applications with dataset distillation are shown in Section 7, and finally we discuss challenges and future directions in Section 8.
2 Background
Before introducing dataset distillation, we first define some notations used in this paper. For a dataset $\mathcal{T}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$, $\bm{x}_{i}\in\mathbb{R}^{d}$, $d$ is the dimension of the input data, and $y_{i}$ is the label. We assume that $(\bm{x}_{i},y_{i})$ for $1\leq i\leq m$ are independent and identically distributed (i.i.d.) random variables drawn from the data generating distribution $\mathcal{D}$. We employ $f_{\bm{\theta}}$ to denote the neural network parameterized by $\bm{\theta}$, and $f_{\bm{\theta}}(\bm{x})$ is the prediction or output of $f_{\bm{\theta}}$ at $\bm{x}$. Moreover, we also define the loss between the prediction and ground truth as $\ell(f_{\bm{\theta}}(\bm{x}),y)$, and the expected risk in terms of $\bm{\theta}$ is defined as
$\mathcal{R}_{\mathcal{D}}(\bm{\theta})=\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}}\left[\ell\left(f_{\bm{\theta}}\left(\bm{x}\right),y\right)\right].$ (1)
Since the data generating distribution $\mathcal{D}$ is unknown, evaluating the expected risk $\mathcal{R}_{\mathcal{D}}$ is impractical. Therefore, a practical way to estimate the expected risk is by the empirical risk $\mathcal{R}_{\mathcal{T}}$, which is defined as
$\mathcal{R}_{\mathcal{T}}(\bm{\theta})=\mathbb{E}_{(\bm{x},y)\sim\mathcal{T}}\left[\ell\left(f_{\bm{\theta}}\left(\bm{x}\right),y\right)\right]=\frac{1}{m}\sum_{i=1}^{m}\ell\left(f_{\bm{\theta}}\left(\bm{x}_{i}\right),y_{i}\right).$ (2)
For the training algorithm alg, $\texttt{alg}(\mathcal{T},\bm{\theta}^{(0)})$ denote the learned parameters returned by empirical risk minimization (ERM) on the dataset $\mathcal{T}$ with the initialized parameter $\bm{\theta}^{(0)}$:
$\texttt{alg}(\mathcal{T},\bm{\theta}^{(0)})=\arg\min_{\bm{\theta}}\mathcal{R}_{\mathcal{T}}(\bm{\theta}).$ (3)
In the deep learning paradigm, gradient descent is the dominant training algorithm to train a neural network by minimizing the empirical risk step by step. Specifically, the network’s parameters are initialized with $\bm{\theta}^{(0)}$, and then the parameter is iteratively updated according to the gradient of empirical risk:
$\bm{\theta}^{(k+1)}=\bm{\theta}^{(k)}-\eta\bm{g}_{\mathcal{T}}^{(k)},$ (4)
where $\eta$ is the learning rate and $\bm{g}_{\mathcal{T}}^{(k)}=\nabla_{\bm{\theta}^{(k)}}\mathcal{R}_{\mathcal{T}}(\bm{\theta}^{(k)})$ is the gradient. We omit $\bm{\theta}^{(0)}$ and use $\texttt{alg}(\mathcal{T})$ if there is no ambiguity for the sake of simplicity. Then, the model $f$ trained on the dataset $\mathcal{T}$ can be denoted as $f_{\texttt{alg}(\mathcal{T})}$.
Because deep learning models are commonly extremely overparameterized, i.e., the number of model parameters overwhelms the number of training examples, the empirical risk readily reaches zero. In this case, the generalization error, which measures the difference between the expected risk and the empirical risk, can be solely equal to the expected risk, which is reflected by test loss or test error in practical pipelines.
2.1 Formalizing Dataset Distillation
Given a target dataset (source training dataset) $\mathcal{T}=\{(\bm{x}_{i},y_{i})\}_{i=1}^{m}$, the objective of dataset distillation is to extract the knowledge of $\mathcal{T}$ into a small synthetic dataset $\mathcal{S}=\{(\bm{s}_{j},y_{j})\}_{j=1}^{n}$, where $n\ll m$, and the model trained on the small distilled dataset $\mathcal{S}$ can achieve a comparable generalization performance to the large original dataset $\mathcal{T}$:
$\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}\atop\bm{\theta}^{(0)}\sim\mathbf{P}}\left[\ell\left(f_{\texttt{alg}(\mathcal{T})}\left(\bm{x}\right),y\right)\right]\simeq\mathbb{E}_{(\bm{x},y)\sim\mathcal{D}\atop\bm{\theta}^{(0)}\sim\mathbf{P}}\left[\ell\left(f_{\texttt{alg}(\mathcal{S})}\left(\bm{x}\right),y\right)\right].$ (5)
Because the training algorithm $\texttt{alg}\left(\mathcal{S},\bm{\theta}^{(0)}\right)$ is determined by the training set $\mathcal{S}$ and the initialized network parameter $\bm{\theta}^{(0)}$, many dataset distillation algorithms will take expectation on $\mathcal{S}$ w.r.t. $\bm{\theta}^{(0)}$ to improve the robustness of the distilled dataset $\mathcal{S}$ to different parameter initialization, and the objective function has the form of $\mathbb{E}_{\bm{\theta}^{(0)}\sim\mathbf{P}}\left[\mathcal{L}\left(\mathcal{S}\right)\right]$. In the following narrative, we omit this expectation w.r.t. initialization $\bm{\theta}^{(0)}$ for the sake of simplicity.
Figure 4: Example synthetic images distilled from CIFAR-10/100 and Tiny ImageNet by matching the training trajectory [13].
3 Meta-Learning Framework
By assuming $\mathcal{R}_{\mathcal{D}}\left({\texttt{alg}(\mathcal{S})}\right)\simeq\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)$, the target dataset $\mathcal{T}$ is employed as the validation set in terms of the model $\texttt{alg}(\mathcal{S})$}. In consequent, the objective of DD can be converted to minimize $\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)$ to improve the generalization performance of $\texttt{alg}(\mathcal{S})$} in terms of $\mathcal{D}$. To this end, dataset distillation falls into the meta-learning area: the hyperparameter $\mathcal{S}$ is updated by the meta (or outer) algorithm, and the base (or inner) algorithm solves the conventional supervised learning problem w.r.t. the synthetic dataset $\mathcal{S}$ [43]; and the dataset distillation task can be formulated to a bilevel optimization problem as follows:
$\mathcal{S}^{\ast}=\arg\min_{\mathcal{S}}\mathcal{R}_{\mathcal{T}}\left({\texttt{alg}(\mathcal{S})}\right)\quad\text{(outer loop)}$ (6)
subject to
$\texttt{alg}(\mathcal{S})=\arg\min_{\bm{\theta}}\mathcal{R}_{\mathcal{S}}\left({\bm{\theta}}\right)\quad\text{(inner loop)}$ (7)
The inner loop optimizes the model parameters based on the synthetic dataset and is often realized by gradient descent for neural networks or regression for kernel method. During the outer loop iteration, the synthetic set is updated by minimizing the model’s risk in terms of the target dataset. With the nested loop, the synthetic dataset gradually converges to one of the optima.
According to the model and optimization methods used in the inner lop, the meta-learning framework of DD can be further classified into two sub-categories of backpropagation through time time (BPTT) approach and kernel ridge regression (KRR) approach.
3.1 Backpropagation Through Time Approach
As shown in the above formulation of bilevel optimization, the objective function of DD can be directly defined as the meta-loss of
$\mathcal{L}(\mathcal{S})=\mathcal{R}_{\mathcal{T}}\left(\texttt{alg}\left(\mathcal{S}\right)\right).$ (8)
Then the distilled dataset is updated by $\mathcal{S}=\mathcal{S}-\alpha\nabla_{\mathcal{S}}\mathcal{L}(\mathcal{S})$ with the step size $\alpha$. However, in the inner loop of Eq. 7, the neural network is trained by iterative gradient descent (Eq. 4), which yields a series of intermediate parameter states of $\{\bm{\mathcal{\theta}}^{(0)},\bm{\mathcal{\theta}}^{(1)},\cdots,\bm{\mathcal{\theta}}^{(T)}\}$, backpropagation through time [115] is required to recursively compute the meta-gradient $\nabla_{\mathcal{S}}\mathcal{L}(\mathcal{S})$:
$\nabla_{\mathcal{S}}\mathcal{L}(\mathcal{S})=\frac{\partial\mathcal{L}}{\partial\mathcal{S}}=\frac{\partial\mathcal{L}}{\partial\bm{\theta}^{(T)}}\left(\sum_{k=0}^{k=T}\frac{\partial\bm{\theta}^{(T)}}{\partial\bm{\theta}^{(k)}}\cdot\frac{\partial\bm{\theta}^{(k)}}{\partial\mathcal{S}}\right),\qquad\frac{\partial\bm{\theta}^{(T)}}{\partial\bm{\theta}^{(k)}}=\prod_{i=k+1}^{T}\frac{\partial\bm{\theta}^{(i)}}{\partial\bm{\theta}^{(i-1)}}.$ (9)
We present the meta-gradient backpropagation of BPTT in Figure 5. Due to the requirement of unrolling the recursive computation graph, BPTT is both computationally expensive and memory demanding, which also severely affects the final distillation performance.
To alleviate the inefficiency in unrolling the long parameter path of $\{\bm{\mathcal{\theta}}^{(0)},\cdots,\bm{\mathcal{\theta}}^{(T)}\}$, Wang et al. [114] adopted a single-step optimization w.r.t. the model parameter from $\bm{\theta}^{(0)}$ to $\bm{\theta}^{(1)}$, and the meta-loss was computed based on $\bm{\theta}^{(1)}$ and the target dataset $\mathcal{T}$:
$\bm{\theta}^{(1)}=\bm{\theta}^{(0)}-\eta\nabla_{\bm{\theta}^{(0)}}\mathcal{R}_{\mathcal{S}}(\bm{\theta}^{(0)})\quad\text{and}\quad\mathcal{L}=\mathcal{R}_{\mathcal{T}}(\bm{\theta}^{(1)}).$ (10)
Therefore, the distilled data $\mathcal{S}$ and learning rate $\eta$ can be efficiently updated via the short-range BPTT as follows:
$\mathcal{S}=\mathcal{S}-\alpha_{\bm{s}_{i}}\nabla\mathcal{L}\quad\text{and}\quad\eta=\eta-\alpha_{\eta}\nabla\mathcal{L}.$ (11)
Unlike freezing distilled labels, Sucholutsky and Schonlau [106] extended the work of Wang et al. [114] by learning a soft-label in the synthetic dataset $\mathcal{S}$; i.e., the label $y$ in the synthetic dataset is also trainable for better information compression, i.e., $y_{i}=y_{i}-\alpha\nabla_{y_{i}}\mathcal{L}$ for $(\bm{s}_{i},y_{i})\in\mathcal{S}$. Similarly, Bohdal et al. [6] also extended the standard example distillation to label distillation by solely optimizing the labels of synthetic datasets. Moreover, these researchers provided improvements on the efficiency of long inner loop optimization via (1) iteratively updating the model parameters $\bm{\theta}$ and the distilled labels $y$, i.e., one outer step followed by only one inner step for faster convergence; and (2) fixing the feature extractor of neural networks and solely updating the last linear layer with ridge regression to avoid second-order meta-gradient computation. Although the BPTT framework has been shown to underperform other algorithms, Deng and Russakovsky [26] empirically demonstrated that adding momentum term and longer unrolled trajectory ($200$ steps) in the inner loop optimization can considerably enhance the distillation performance; and the inner loop of model training becomes
$\bm{\theta}^{(k+1)}=\bm{\theta}^{(k)}-\eta\bm{m}^{(k+1)},$ (12)
where
$\bm{m}^{(k+1)}=\beta\bm{m}^{(k)}+\nabla_{\bm{\theta}^{(k)}}\mathcal{R}_{\mathcal{T}}(\bm{\theta}^{(k)})\quad\text{s.t.}\quad\bm{m}^{(0)}=\bm{0}.$ (13)
Figure 5: (a) Meta-gradient backpropagation in BPTT [132]; (b) Meta-gradient backpropagation in kernel ridge regression; and (c) Meta-gradient backpropagation in gradient matching.
3.2 Kernel Ridge Regression Approach
Although multistep gradient descent can gradually approach the optimal network parameters in terms of the synthetic dataset during the inner loop, this iterative algorithm makes the meta-gradient backpropagation highly inefficient, as shown in BPTT. Considering the existence of closed-form solutions in the kernel regression regime, Nguyen et al. [86] replaced the neural network in the inner loop with a kernel model, which bypasses the recursive backpropagation of the meta-gradient. For the regression model $f(\bm{x})=\bm{w}^{\top}\psi(\bm{x})$, where $\psi(\cdot)$ is a nonlinear mapping and the corresponding kernel is $K(\bm{x},\bm{x}^{\prime})=\langle\psi(\bm{x}),\psi(\bm{x}^{\prime})\rangle$, there exists a closed-form solution for $\bm{w}$ when the regression model is trained on $\mathcal{S}$ with kernel ridge regression (KRR):
$\bm{w}=\psi(X_{s})^{\top}\left(\mathbf{K}_{X_{s}X_{s}}+\lambda I\right)^{-1}y_{s},$ (14)
where $\mathbf{K}_{X_{s}X_{s}}=[K(\bm{s}_{i},\bm{s}_{j})]_{ij}\in\mathbb{R}^{n\times n}$ is called the kernel matrix or Gram matrix associated with $K$ and the dataset $\mathcal{S}$, and $\lambda>0$ is a fixed regularization parameter [89]. Therefore, the mean square error (MSE) of predicting $\mathcal{T}$ with the model trained on $\mathcal{S}$ is
$\mathcal{L}(\mathcal{S})=\frac{1}{2}\left\|y_{t}-\mathbf{K}_{X_{t}X_{s}}\left(\mathbf{K}_{X_{s}X_{s}}+\lambda I\right)^{-1}y_{s}\right\|^{2},$ (15)
where $\mathbf{K}_{X_{t}X_{s}}=[K(\bm{x}_{i},\bm{s}_{j})]_{ij}\in\mathbb{R}^{m\times n}$. Then the distilled dataset is updated via the meta-gradient of the above loss. Due to the closed-form solution in KRR, $\bm{\theta}$ does not require an iterative update and the backward pass of the gradient thus bypasses the recursive computation graph, as shown in Figure 5.
In the KRR regime, the synthetic dataset $\mathcal{S}$ can be directly updated by backpropagating the meta-gradient through the kernel function. Although this formulation is solid, this algorithm is designed in the KRR scenario and only employs simple kernels, which causes performance drops when the distilled dataset is transferred to train neural networks. Jacot et al. [47] proposed the neural tangent kernel (NTK) theory that proves the equivalence between training infinite-width neural networks and kernel regression. With this equivalence, Nguyen et al. [87] employed infinite-width networks as the kernel for dataset distillation, which narrows the gap between the scenarios of KRR and deep learning.
However, every entry in the kernel matrix must be calculated separately via the kernel function, and thus computing the kernel matrix $\mathbf{K}_{X_{s}X_{t}}$ has the time complexity of $\mathcal{O}(|\mathcal{T}||\mathcal{S}|)$, which is severely inefficient for large-scale datasets with huge $|\mathcal{T}|$. To tackle this problem, Loo et al. [74] replaced the NTK kernel with neural network Gaussian process (NNGP) kernel that only considers the training dynamic of the last-layer classifier for speed up. With this replacement, the random features $\psi(\bm{x})$ and $\psi(\bm{s})$ can be explicitly computed via multiple sampling from the Gaussian process, and thus the kernel matrix computation can be decomposed into random feature calculation and random feature matrix multiplication. Because matrix multiplication requires negligible amounts of time for small distilled datasets, the time complexity of kernel matrix computation degrades to $\mathcal{O}(|\mathcal{T}|+|\mathcal{S}|)$. In addition, these authors demonstrated two issues of MSE loss (Eq. 15) used in [86] and [87]: (1) over-influence on corrected data: the correctly classified examples in $\mathcal{T}$ can induce larger loss than the incorrectly classified examples; and (2) unclear probabilistic interpretation for classification tasks. To overcome these issues, they propose to apply a cross-entropy (CE) loss with Platt scaling [91] to replace the MSE loss:
$\mathcal{L}_{\tau}=\text{CE}(y_{t},\hat{y}_{t}/\tau),$ (16)
where $\tau$ is a positive learned temperature scaling parameter, and $\hat{y}_{t}=\mathbf{K}_{X_{t}X_{s}}\left(\mathbf{K}_{X_{s}X_{s}}+\lambda I\right)^{-1}y_{s}$ is still calculated using the KRR formula.
A similar efficient method was also proposed by Zhou et al. [132], which also focused on solving the last-layer in neural networks with KRR. Specifically, the neural network $f_{\bm{\theta}}=g_{\bm{\theta}_{2}}\circ h_{\bm{\theta}_{1}}$ can be decomposed with the feature extractor $h$ and the linear classifier $g$. Then these coworkers fixed the feature extractor $h$ and the linear classifier $g$ possesses a closed-form solution with KRR; and the the distilled data are accordingly optimized with the MSE loss:
$\mathcal{L}(\mathcal{S})=\frac{1}{2}\left\|y_{t}-\mathbf{K}_{X_{t}X_{s}}^{\bm{\theta}_{1}^{(k)}}\left(\mathbf{K}_{X_{s}X_{s}}^{\bm{\theta}_{1}^{(k)}}+\lambda I\right)^{-1}y_{s}\right\|^{2}$ (17)
and
$\bm{\theta}_{1}^{(k)}=\bm{\theta}_{1}^{(k-1)}-\eta\nabla_{\bm{\theta}_{1}^{(k-1)}}\mathcal{R}_{\mathcal{S}}(\bm{\theta}_{1}^{(k-1)}),$ (18)
where the kernel matrices $\mathbf{K}_{X_{t}X_{s}}^{\bm{\theta}_{1}}$ and $\mathbf{K}_{X_{s}X_{s}}^{\bm{\theta}_{1}}$ are induced by $h_{\bm{\theta}_{1}}(\cdot)$. Notably, although the feature extractor $h_{\bm{\theta}_{1}}$ is continuously updated, the classifier $g_{\bm{\theta}_{2}}$ is directly solved by KRR, and thus the meta-gradient backpropagation side-steps the recursive computation graph.
3.3 Discussion
From the loss surface perspective [68], the meta-learning framework of minimizing $\mathcal{R}_{\mathcal{T}}\left(\texttt{alg}\left(\mathcal{S}\right)\right)$ can be considered to mimic the local minima of target data with the distilled data. However, the loss landscape w.r.t. parameters is closely related to the network architecture, while only one type of small network is used in the BPTT approach. Consequently, there is a moderate performance drop when the distilled dataset is employed to train other complicated networks. Moreover, a long unrolled trajectory and second-order gradient computation are also two key challenges for BPTT approach, which hinder its efficiency. The KRR approach compensates for these shortcomings by replacing networks with the nonparametric kernel model which admits closed-form solution. Although KRR is nonparametric and does not involve neural networks during the distillation process, previous research has shown that the training dynamic of neural networks is equal to the kernel method when the width of networks becomes infinite [47, 35, 5, 5], which partially guarantees the feasibility of the kernel regression approach and explains its decent performance when transferred to wide neural networks.
4 Data Matching Framework
Although it is not feasible to explicitly extract information from target data and then inject them into synthetic data, information distillation can be achieved by implicitly aligning the byproducts of target data and synthetic data from different aspects; and this byproduct matching allows synthetic data to imitate the influence of target data on model training. The objective function of data matching can be summarized as follows.
$\mathcal{L}(\mathcal{S})=\sum_{k=0}^{T}D\left(\phi(\mathcal{S},\bm{\theta}^{(k)}),\phi(\mathcal{T},\bm{\theta}^{(k)})\right)$ (19)
subject to
$\bm{\theta}^{(k)}=\bm{\theta}^{(k-1)}-\eta\nabla_{\bm{\theta}^{(k-1)}}\mathcal{R}_{\mathcal{S}}\left(\bm{\theta}^{(k-1)}\right),$ (20)
where $D(\cdot,\cdot)$ is a distance function, and $\phi(\cdot)$ maps the dataset $\mathcal{S}$ or $\mathcal{T}$ to other informative spaces, such as gradient, parameter, and feature spaces. In practical implementation, the full datasets of $\mathcal{S}$ and $\mathcal{T}$ are often replaced with random sampled batches of $\mathcal{B}_{\mathcal{S}}$ and $\mathcal{B}_{\mathcal{T}}$ for memory saving and faster convergence.
Compared to the aforementioned meta-learning framework, the data matching loss not only focuses on the final parameter $\texttt{alg}(\mathcal{S})$ but also supervises the intermediate states, as shown in the sum operation $\sum_{k=0}^{T}$. By this, the distilled data can better imitate the influence of target data on training networks at different training stages.
4.1 Gradient Matching Approach
To achieve comparable generalization performance, a natural approach is to imitate the model parameters, i.e., matching the training trajectories introduced by $\mathcal{S}$ and $\mathcal{T}$. With a fixed parameter initialization, the training trajectory of $\{\bm{\mathcal{\theta}}^{(0)},\bm{\mathcal{\theta}}^{(1)},\cdots,\bm{\mathcal{\theta}}^{(T)}\}$ is equal to a series of gradients $\{\bm{g}^{(0)},\cdots,\bm{g}^{(T)}\}$. Therefore, matching the gradients induced by $\mathcal{S}$ and $\mathcal{T}$ is a convincing proxy to mimic the influence on model parameters [129], and the objective function is formulated as
$\mathcal{L}(\mathcal{S})=\sum_{k=0}^{T-1}D\left(\nabla_{\bm{\theta}^{(k)}}\mathcal{R}_{\mathcal{S}}\left(\bm{\theta}^{(k)}\right),\nabla_{\bm{\theta}^{(k)}}\mathcal{R}_{\mathcal{T}}\left(\bm{\theta}^{(k)}\right)\right).$ (21)
The difference $D$ between gradients is measured in the layer-wise aspect:
$D\left(\bm{g}_{\mathcal{S}},\bm{g}_{\mathcal{T}}\right)=\sum_{l=1}^{L}\texttt{dis}\left(\bm{g}_{\mathcal{S}}^{l},\bm{g}_{\mathcal{T}}^{l}\right),$ (22)
where $\bm{g}^{l}$ denotes the gradient of $i$-th layer, and dis is the sum of cosine distance as follows:
$\texttt{dis}\left(\bm{A},\bm{B}\right)=\sum_{i=1}^{\texttt{out}}\left(1-\frac{\bm{A}_{i}\bm{B}_{i}}{\|\bm{A}_{i}\|\|\bm{B}_{i}\|}\right),$ (23)
where out denotes the number of output channels for specific layer; and $\bm{A}_{i}$ and $\bm{B}_{i}$ are the flatten gradients in the $i$-th channel.
To improve the convergence speed in practical implementation, Zhao et al. [129] proposed to match gradient in class-wise as follows:
$\mathcal{L}^{(k)}=\sum_{c=1}^{C}D\left(\nabla_{\bm{\theta}^{(k)}}\mathcal{R}\left(\mathcal{B}_{\mathcal{S}}^{c},\bm{\theta}^{(k)}\right),\nabla_{\bm{\theta}^{(k)}}\mathcal{R}\left(\mathcal{B}_{\mathcal{T}}^{c},\bm{\theta}^{(k)}\right)\right),$ (24)
where $c$ is the class index and $\mathcal{B}_{\mathcal{S}}^{c}$ and $\mathcal{B}_{\mathcal{T}}^{c}$ denotes the batch examples belong to the $c$-th class. However, according to Lee et al. [63], the class-wise gradient matching pays much attention to the class-common features and overlooks the class-discriminative features in the target dataset, and the distilled synthetic dataset $\mathcal{S}$ does not possess enough class-discriminative information, especially when the target dataset is fine-grained; i.e., class-common features are the dominant. Based on this finding, these coworkers proposed an improved objective function of
$\mathcal{L}^{(k)}=D\left(\sum_{c=1}^{C}\nabla_{\bm{\theta}^{(k)}}\mathcal{R}\left(\mathcal{B}_{\mathcal{S}}^{c},\bm{\theta}^{(k)}\right),\sum_{c=1}^{C}\nabla_{\bm{\theta}^{(k)}}\mathcal{R}\left(\mathcal{B}_{\mathcal{T}}^{c},\bm{\theta}^{(k)}\right)\right)$ (25)
to better capture constractive signals between different classes through the sum of loss gradients between classes. A similar approach was proposed by Jiang et al. [48], which employed both intraclass and interclass gradient matching by combining Eq. 24 and Eq. 25. Moreover, these researchers measure the difference of gradients by considering the magnitude instead of only considering the angle, i.e., the cosine distance, and the distance function of Eq. 23 is improved to
$\texttt{dis}\left(\bm{A},\bm{B}\right)=\sum_{i=1}^{\texttt{out}}\left(1-\frac{\bm{A}_{i}\bm{B}_{i}}{\|\bm{A}_{i}\|\|\bm{B}_{i}\|}+\|\bm{A}_{i}-\bm{B}_{i}\|\right).$ (26)
To alleviate easy overfitting on the small dataset $\mathcal{S}$, Kim et al. [54] proposed to perform inner loop optimization on the target dataset $\mathcal{T}$ instead of $\mathcal{S}$, i.e., replaced the parameter update of Eq. 20 with
$\bm{\theta}^{(k+1)}=\bm{\theta}^{(k)}-\eta\nabla_{\bm{\theta}^{(k)}}\mathcal{R}_{\mathcal{T}}\left(\bm{\theta}^{(k)}\right).$ (27)
Although data augmentation facilitates a large performance increase in conventional network training, conducting augmentation on distilled datasets demonstrates no improvement on the final test accuracy, because the characteristics of synthetic images are different from those of natural images and are not optimized under the supervision of various transformations. To leverage data augmentation on synthetic datasets, Zhao and Bilen [126] designed data Siamese augmentation (DSA) that homologously augments the distilled data and the target data during the distillation process as follows:
$\mathcal{L}=D\left(\nabla_{\bm{\theta}}\mathcal{R}\left(\mathcal{A}_{\omega}\left(\mathcal{B}_{\mathcal{S}}\right),\bm{\theta}\right),\nabla_{\bm{\theta}}\mathcal{R}\left(\mathcal{A}_{\omega}\left(\mathcal{B}_{\mathcal{T}}\right),\bm{\theta}\right)\right),$ (28)
where $\mathcal{A}$ is a family of image transformations such as cropping and flipping that are parameterized with $\omega^{\mathcal{S}}$ and $\omega^{\mathcal{T}}$ for synthetic and target data, respectively. In DSA, the augmented form of distilled data has a consistent correspondence w.r.t. the augmented form of the target data, i.e., $\omega^{\mathcal{S}}=\omega^{\mathcal{T}}=\omega$; and $\omega$ is randomly picked from $\mathcal{A}$ at different iterations. Notably, the transformation $\mathcal{A}$ requires to be differentiable w.r.t. the synthetic dataset $\mathcal{S}$ for backpropagation:
$\frac{\partial D(\cdot)}{\partial\mathcal{S}}=\frac{\partial D(\cdot)}{\partial\nabla_{\theta}\mathcal{L}(\cdot)}\frac{\partial\nabla_{\theta}\mathcal{L}(\cdot)}{\partial\mathcal{A}(\cdot)}\frac{\partial\mathcal{A}(\cdot)}{\partial\mathcal{S}}.$ (29)
Through setting $\omega^{\mathcal{S}}=\omega^{\mathcal{T}}$, DSA permits the knowledge transfer from the transformation of target images to the corresponding transformation of synthetic images. Consequently, the augmented synthetic images also possess meaningful characteristics of the natural images. Due to its superior compatibility, DSA has been widely used in many data matching methods.
(a) Trajectory matching
(b) Distribution matching
Figure 6: (a) An illustration of trajectory matching [13]; (b) Compared to gradient matching approach, distribution matching approach can more comprehensively cover the data distribution in feature space.
4.2 Trajectory Matching Approach
Unlike circuitously matching the gradients, Cazenavette et al. [13] directly matched the long-range training trajectory between the target dataset and the synthetic dataset. Specifically, they train models on the target dataset $\mathcal{T}$ and collect the expert training trajectory into a buffer in advance. Then the ingredients in the buffer are randomly selected to initialize the networks for training $\mathcal{S}$. After collecting the trajectories of $\mathcal{S}$, the synthetic dataset is updated by matching their parameters, as shown in Figure 6(a). The objective loss of trajectory matching is defined as
$\mathcal{L}=\frac{\|\bm{\theta}^{(k+N)}-\bm{\theta}_{\mathcal{T}}^{(k+M)}\|_{2}^{2}}{\|\bm{\theta}_{\mathcal{T}}^{(k)}-\bm{\theta}_{\mathcal{T}}^{(k+M)}\|_{2}^{2}},$ (30)
where $\bm{\theta}_{\mathcal{T}}$ denotes the target parameter by training the model on $\mathcal{T}$, which is stored in the buffer, and $\bm{\theta}^{(k+N)}$ are the parameter obtained by training the model on $\mathcal{S}$ for $N$ epochs with the initialization of $\bm{\theta}_{\mathcal{T}}^{(k)}$. The denominator in the loss function is for normalization.
Although trajectory matching demonstrated empirical success, Du et al. [29] argued that $\bm{\theta}^{(k)}$ is generally induced by training on $\mathcal{S}$ for $k$ epochs during natural learning process, while it was directly initialized with $\bm{\theta}_{\mathcal{T}}^{(k)}$ in trajectory matching, which causes the accumulated trajectory error that measures the difference between the parameters from real trajectories of $\mathcal{S}$ and $\mathcal{T}$. Because the accumulated trajectory error is induced by the mismatch between $\bm{\theta}_{\mathcal{T}}^{(k)}$ and real $\bm{\theta}^{(k)}$ from the natural training, these researchers alleviate this error by adding random noise to when initialize $\bm{\theta}^{(k)}$ to improve robustness w.r.t. $\bm{\theta}_{\mathcal{T}}^{(k)}$.
Compared to matching gradients, while trajectory matching side-steps second-order gradient computation, unfortunately, unrolling $N$ SGD updates are required during meta-gradient backpropagation due to the existence of $\bm{\theta}^{(k+N)}$ in Eq. 30. The unrolled gradient computation significantly increases the memory burden and impedes scalability. By disentangling the meta-gradient w.r.t. synthetic examples into two passes, Cui et al. [25] greatly reduced the memory required by trajectory matching and successfully scaled trajectory matching approach to the large ImageNet-1K dataset [95]. Motivated by knowledge distillation [38], these scholars also proposed assigning soft-labels to synthetic examples with pretrained models in the buffer; and the soft-labels help learn intraclass information and consequently improve distillation performance.
4.3 Distribution Matching Approach
Figure 7: An illustration of distribution matching. $\psi$ is the feature extractor, and red dashed lines denote the backpropagation of gradients.
Although the above parameter-wise matching shows satisfying performance, Zhao and Bilen [128] visualized the distilled data in two dimension and revealed that there is a large distribution discrepancy between the distilled data and the target data. In other words, the distilled dataset cannot comprehensively cover the data distribution in feature space, as shown in Figure 6(b). Based on this discovery, these researchers proposed to match the synthetic and target data from the distribution perspective for dataset distillation. Specifically, they employed the pretrained feature extractor $\psi_{\bm{v}}$ with the parameter $\bm{v}$ to achieve mapping from the input space to the feature space. The synthetic data are optimized according to the objective function
$\mathcal{L}(\mathcal{S})=\sum_{c=1}^{C}\|\psi_{\bm{v}}(\mathcal{B}_{\mathcal{S}}^{c})-\psi_{\bm{v}}(\mathcal{B}_{\mathcal{T}}^{c})\|^{2},$ (31)
where $c$ denotes the class index. An illustration of distribution matching is also presented Figure 7. As shown in Eq. 31, distribution matching does not rely on the model parameters and drops bilevel optimization for less memory requirement, whereas it empirically underperforms the above gradient and trajectory matching approaches.
Wang et al. [113] improved the distribution matching from several aspects: (1) using multiple-layer features other than only the last-layer features for matching, and the feature matching loss is formulated as follows:
$\mathcal{L}_{\text{f}}=\sum_{c=1}^{C}\sum_{l=1}^{L}|f_{\bm{\theta}}^{l}(\mathcal{B}_{\mathcal{S}}^{c})-f_{\bm{\theta}}^{l}(\mathcal{T}_{\mathcal{S}}^{c})|^{2},$ (32)
where $f_{\bm{\theta}}^{l}$ denotes the $l$-th layer features and $L$ is the total number of network layers; (2) recalling the bilevel optimization that updates $\mathcal{S}$ with different model parameters by inserting the inner loop of $\bm{\theta}^{(k+1)}=\bm{\theta}^{(k)}-\eta\nabla_{\bm{\theta}^{(k)}}\mathcal{R}_{\mathcal{S}}(\bm{\theta}^{(k)})$; and (3) proposing the discrimination loss in the last-layer feature space to enlarge the class distinction of synthetic data. Specifically, the synthetic feature center $\bar{f}^{\mathcal{S}}_{c,L}$ of each category $c$ is obtained by averaging the batch $\mathcal{B}_{\mathcal{S}}^{c}$. The objective of discrimination loss is to improve the classification ability of the feature center $\bar{\bm{F}}_{L}^{\mathcal{S}}=[\bar{f}^{\mathcal{S}}_{1,L},\bar{f}^{\mathcal{S}}_{2,L},\cdots,\bar{f}^{\mathcal{S}}_{C,L}]$ on predicting the real data feature $\bm{F}_{L}^{\mathcal{T}}=[{f}^{\mathcal{T}}_{1,L},{f}^{\mathcal{T}}_{1,L},\cdots,{f}^{\mathcal{T}}_{C,L}]$. The logits are first calculated by
$\mathbf{O}=\left\langle\mathbf{\bm{F}}_{L}^{\mathcal{T}},\left(\bar{\bm{F}}_{L}^{\mathcal{S}}\right)^{\top}\right\rangle,$ (33)
where $\mathbf{O}\in\mathbb{R}^{N\times C}$ contains the logits of $N=C\times|\mathcal{B}_{\mathcal{T}}^{c}|$ target data points. Then the probability $p_{i}$ in terms of the ground-truth label is derived via $p_{i}=\text{softmax}(\mathbf{O}_{i})$; and the classification loss is
$\mathcal{L}_{\text{d}}=\frac{1}{N}\sum_{i=1}^{N}\log p_{i}.$ (34)
Therefore, the total loss in [113] for dataset distillation is $\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{f}}+\beta\mathcal{L}_{\text{d}}$, where $\beta$ is a positive factor for balancing the feature matching and discrimination loss.
4.4 Discussion
The gradient matching approach can be considered as to be short-range parameter matching, while its backpropagation requires second-order gradient computation. Although the trajectory matching approach make ups for this imperfection, the long-range trajectory introduces a recursive computation graph during meta-gradient backpropagation. Different from matching in parameter space, distribution matching approach employs the feature space as the match proxy and also bypasses second-order gradient computation. Although distribution matching has advantages in scalability w.r.t. high-dimensional data, it empirically underperforms trajectory matching, which might be attributed to the mismatch between the comprehensive distribution and decent distillation performance. In detail, distribution matching achieves DD by mimicking features of the target data, so that the distilled data are evenly distributed in the feature space. However, not all features are equally important and distribution matching might waste the budget on imitating less informative features, thereby undermining the distillation performance.
5 Factorized Dataset Distillation
In representation learning, although the image data are in the extremely high-dimensional space, they may lie on a low-dimensional manifold and rely on a few features, and one can recover the source image from the low-dimensional features with specific decoders [3, 122]. For this reason, it is plausible to implicitly learn synthetic datasets by optimizing their factorized features and corresponding decoders, which is termed as factorized dataset distillation, as shown in Figure 8. To be in harmony with decoders, we recall the notion of a feature as code in terms of the dataset. By factorizing synthetic datasets with a combination of codes and decoders, the compression ratio can be further decreased, and information redundancy in distilled images can also be reduced. According to the learnability of the code and decoder, we classify factorized dataset distillation into three categories of code-based DD, decoder-based DD, and code-decoder DD.
(a) Nonfactorized dataset distillation
(b) Factorized dataset distillation
Figure 8: Schematic diagrams of nonfactorized dataset distillation and factorized dataset distillation.
Code-based DD aims to learn a series of low-dimensional codes for generating highly informative images through a specific generator. In [127], the vectors that are put into the GAN generator are learned to producing informative images. Concretely, these investigators inverse real examples with a GAN generator and collect corresponding latent codes. Then, the latent vectors are further optimized with the distribution matching algorithm. By this, the optimized latent vectors can induce more informative synthetic examples with the pretrained GAN generator. In addition to use the GAN, Kim et al. [54] employed a deterministic multi-formation function $\texttt{multi-form}(\cdot)$ as the decoder to create synthetic data from fewer condensed data, i.e., the synthetic data are generated by $\mathcal{S}=\texttt{multi-form}(\mathcal{C})$. Then the condensed data $\mathcal{C}$ is optimized in an end-to-end fashion by the gradient matching approach.
Different from code-based DD, decoder-based DD solely learns decoders that are used to produce highly informative data. In [105], a generative teaching network (GTN) that employs a trainable network was proposed to generate synthetic images from random noise based on given labels, and the meta-gradients are backpropagated via BPTT to update the GTN rather than the synthetic data.
Naturally, code-decoder DD combines code-based and decoder-based DD and allows the training on both codes and decoders. Deng and Russakovsky [26] generated synthetic images via the matrix multiplication between codes and decodes, which they call the memory and addressing function. Specifically, they use the memory $\mathcal{M}=\{\bm{b}_{1},\cdots,\bm{b}_{K}\}$ to store the bases $\bm{b}_{i}\in\mathbb{R}^{d}$ that have the same dimension as target data; and $r$ addressing functions of $\mathcal{A}=\{\bm{A}_{1},\cdots,\bm{A}_{r}\}$ are used to recover the synthetic images according to the given one-hot label $\bm{y}$ as follows:
$\bm{s}_{i}^{\top}=\bm{y}^{T}\bm{A}_{i}[\bm{b}_{1},\cdots,\bm{b}_{K}]^{\top},$ (35)
where the one-hot label $\bm{y}\in\mathbb{R}^{C\times 1}$, and $\bm{A}_{i}\in\mathbb{R}^{C\times K}$, and $C$ is the number of classes. By plugging different addressing function $\bm{A}_{i}$ into Eq. 35, a total number of $r$ synthetic images can be generated from the bases for each category. During the distillation process, these two elements of $\mathcal{M}$ and $\mathcal{A}$ were optimized with the meta-learning framework. Different from matrix multiplication in [26], Liu et al. [71] employ hallucinator networks as decoders to generate synthetic data from basis. Specifically, a hallucinator network is consist of three parts of a encoder enc, an affine transformation with trainable scale $\sigma$ and shift $\mu$, and a decoder dec, then a basis $\bm{b}$ is put into the hallucinator network to generate corresponding synthetic image $\bm{s}$ as follows:
$\mathbf{f}_{1}=\texttt{enc}(\bm{b}),\quad\mathbf{f}_{2}=\sigma\times\mathbf{f}_{1}+\mu,\quad\bm{s}=\texttt{dec}(\mathbf{f}_{2}),$ (36)
where the multiplication is element-wise operation. In addition, to enlarge the knowledge divergence encoded by different hallucinator networks for efficient compression, these researchers proposed an adversarial contrastive constraint that maximizes the divergence of features of different generated synthetic data. A similar code-decoder factorization method is also presented in Lee et al. [62], where they adopted an improved distribution matching to optimize the latent codes and decoders.
Through implicitly learning the latent codes and decoders, factorized DD possesses advantages such as more compact representation and shared representation across classes that consequently improve the dataset distillation performance. Notably, this code-decoder factorization is compatible with the aforementioned distillation approaches. Therefore, the exploration of synthetic data generation and distillation frameworks can promote dataset distillation in parallel.
6 Performance Comparison
Table 1: Performance comparison of different dataset distillation methods on different datasets. Abbreviations of GM, TM, DM are for gradient matching, trajectory matching, and distribution matching, respectively. Whole denotes the test accuracy in terms of the whole target dataset.
Methods Schemes MNIST FashionMNIST SVHN CIFAR-10 CIFAR-100 Tiny ImageNet
11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$ 11 $10$ $50$
Random -
$64.9$
$\scriptscriptstyle\pm 3.5$
$95.1$
$\scriptscriptstyle\pm 0.9$
$97.9$
$\scriptscriptstyle\pm 0.2$
$51.4$
$\scriptscriptstyle\pm 3.8$
$73.8$
$\scriptscriptstyle\pm 0.7$
$82.5$
$\scriptscriptstyle\pm 0.7$
$14.6$
$\scriptscriptstyle\pm 1.6$
$35.1$
$\scriptscriptstyle\pm 4.1$
$70.9$
$\scriptscriptstyle\pm 0.9$
$14.4$
$\scriptscriptstyle\pm 2.0$
$26.0$
$\scriptscriptstyle\pm 1.2$
$43.4$
$\scriptscriptstyle\pm 1.0$
$4.2$
$\scriptscriptstyle\pm 0.3$
$14.6$
$\scriptscriptstyle\pm 0.5$
$30.0$
$\scriptscriptstyle\pm 0.4$
$1.4$
$\scriptscriptstyle\pm 0.1$
$5.0$
$\scriptscriptstyle\pm 0.2$
$15.0$
$\scriptscriptstyle\pm 0.4$
Herding -
$89.2$
$\scriptscriptstyle\pm 1.6$
$93.7$
$\scriptscriptstyle\pm 0.3$
$94.8$
$\scriptscriptstyle\pm 0.2$
$67.0$
$\scriptscriptstyle\pm 1.9$
$71.1$
$\scriptscriptstyle\pm 0.7$
$71.9$
$\scriptscriptstyle\pm 0.8$
$20.9$
$\scriptscriptstyle\pm 1.3$
$50.5$
$\scriptscriptstyle\pm 3.3$
$72.6$
$\scriptscriptstyle\pm 0.8$
$21.5$
$\scriptscriptstyle\pm 1.2$
$31.6$
$\scriptscriptstyle\pm 0.7$
$40.4$
$\scriptscriptstyle\pm 0.6$
$8.4$
$\scriptscriptstyle\pm 0.3$
$17.3$
$\scriptscriptstyle\pm 0.3$
$33.7$
$\scriptscriptstyle\pm 0.5$
$2.8$
$\scriptscriptstyle\pm 0.2$
$6.3$
$\scriptscriptstyle\pm 0.2$
$16.7$
$\scriptscriptstyle\pm 0.3$
DD [114] BPTT -
$79.5$
$\scriptscriptstyle\pm 8.1$
$36.8$
$\scriptscriptstyle\pm 1.2$
LD [6] BPTT
$60.9$
$\scriptscriptstyle\pm 3.2$
$87.3$
$\scriptscriptstyle\pm 0.7$
$93.3$
$\scriptscriptstyle\pm 0.3$
$25.7$
$\scriptscriptstyle\pm 0.7$
$38.3$
$\scriptscriptstyle\pm 0.4$
$42.5$
$\scriptscriptstyle\pm 0.4$
$11.5$
$\scriptscriptstyle\pm 0.4$
DC [129] GM
$91.7$
$\scriptscriptstyle\pm 0.5$
$94.7$
$\scriptscriptstyle\pm 0.2$
$98.8$
$\scriptscriptstyle\pm 0.2$
$70.5$
$\scriptscriptstyle\pm 0.6$
$82.3$
$\scriptscriptstyle\pm 0.4$
$83.6$
$\scriptscriptstyle\pm 0.4$
$31.2$
$\scriptscriptstyle\pm 1.4$
$76.1$
$\scriptscriptstyle\pm 0.6$
$82.3$
$\scriptscriptstyle\pm 0.3$
$28.3$
$\scriptscriptstyle\pm 0.5$
$44.9$
$\scriptscriptstyle\pm 0.5$
$53.9$
$\scriptscriptstyle\pm 0.5$
$13.9$
$\scriptscriptstyle\pm 0.4$
$32.3$
$\scriptscriptstyle\pm 0.3$
$42.8$
$\scriptscriptstyle\pm 0.4$
DSA [126] GM
$88.7$
$\scriptscriptstyle\pm 0.6$
$97.8$
$\scriptscriptstyle\pm 0.1$
$\mathbf{99.2}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$70.6$
$\scriptscriptstyle\pm 0.6$
$86.6$
$\scriptscriptstyle\pm 0.3$
$88.7$
$\scriptscriptstyle\pm 0.2$
$27.5$
$\scriptscriptstyle\pm 1.4$
$\mathbf{79.2}$ $\scriptscriptstyle\mathbf{\pm 0.5}$
$84.4$
$\scriptscriptstyle\pm 0.4$
$28.8$
$\scriptscriptstyle\pm 0.7$
$52.1$
$\scriptscriptstyle\pm 0.5$
$60.6$
$\scriptscriptstyle\pm 0.5$
$13.9$
$\scriptscriptstyle\pm 0.4$
$32.3$
$\scriptscriptstyle\pm 0.3$
$42.8$
$\scriptscriptstyle\pm 0.4$
DCC [63] GM - - - - - -
$47.5$
$\scriptscriptstyle\pm 2.6$
$\mathbf{80.5}$ $\scriptscriptstyle\mathbf{\pm 0.6}$ $\mathbf{87.2}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
$34.0$
$\scriptscriptstyle\pm 0.7$
$54.5$
$\scriptscriptstyle\pm 0.5$
$64.2$
$\scriptscriptstyle\pm 0.4$
$14.6$
$\scriptscriptstyle\pm 0.3$
$33.5$
$\scriptscriptstyle\pm 0.3$
$39.3$
$\scriptscriptstyle\pm 0.4$
MTT [13] TM
$91.4$
$\scriptscriptstyle\pm 0.9$
$97.3$
$\scriptscriptstyle\pm 0.1$
$98.5$
$\scriptscriptstyle\pm 0.1$
$75.1$
$\scriptscriptstyle\pm 0.9$
$\mathbf{87.2}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
$88.3$
$\scriptscriptstyle\pm 0.1$
$46.3$
$\scriptscriptstyle\pm 0.8$
$65.3$
$\scriptscriptstyle\pm 0.7$
$71.6$
$\scriptscriptstyle\pm 0.2$
$24.3$
$\scriptscriptstyle\pm 0.3$
$40.1$
$\scriptscriptstyle\pm 0.4$
$47.7$
$\scriptscriptstyle\pm 0.2$
$8.8$
$\scriptscriptstyle\pm 0.3$
$23.2$
$\scriptscriptstyle\pm 0.2$
$\mathbf{28.0}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
FTD [29] TM - - - - - - - - -
$46.8$
$\scriptscriptstyle\pm 0.3$
$\mathbf{66.6}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{73.8}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
$25.2$
$\scriptscriptstyle\pm 0.2$
$\mathbf{43.4}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
$50.7$
$\scriptscriptstyle\pm 0.3$
$\mathbf{10.4}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{24.5}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ -
TESLA [25] TM - - - - - - - - -
$48.5$
$\scriptscriptstyle\pm 0.8$
$\mathbf{66.4}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ $\mathbf{72.6}$ $\scriptscriptstyle\mathbf{\pm 0.7}$
$24.8$
$\scriptscriptstyle\pm 0.4$
$41.7$
$\scriptscriptstyle\pm 0.3$
$47.9$
$\scriptscriptstyle\pm 0.3$
$7.7$
$\scriptscriptstyle\pm 0.2$
$18.8$
$\scriptscriptstyle\pm 1.3$
$\mathbf{27.9}$ $\scriptscriptstyle\mathbf{\pm 1.2}$
DM [128] DM
$89.2$
$\scriptscriptstyle\pm 1.6$
$97.3$
$\scriptscriptstyle\pm 0.3$
$94.8$
$\scriptscriptstyle\pm 0.2$
$26.0$
$\scriptscriptstyle\pm 0.8$
$48.9$
$\scriptscriptstyle\pm 0.6$
$63.0$
$\scriptscriptstyle\pm 0.4$
$11.4$
$\scriptscriptstyle\pm 0.3$
$29.7$
$\scriptscriptstyle\pm 0.3$
$43.6$
$\scriptscriptstyle\pm 0.4$
$3.9$
$\scriptscriptstyle\pm 0.2$
$12.9$
$\scriptscriptstyle\pm 0.4$
$24.1$
$\scriptscriptstyle\pm 0.3$
CAFE [113] DM
$90.8$
$\scriptscriptstyle\pm 0.5$
$97.5$
$\scriptscriptstyle\pm 0.1$
$98.9$
$\scriptscriptstyle\pm 0.2$
$73.7$
$\scriptscriptstyle\pm 0.7$
$83.0$
$\scriptscriptstyle\pm 0.3$
$88.2$
$\scriptscriptstyle\pm 0.3$
$42.9$
$\scriptscriptstyle\pm 3.0$
$77.9$
$\scriptscriptstyle\pm 0.6$
$82.3$
$\scriptscriptstyle\pm 0.4$
$31.6$
$\scriptscriptstyle\pm 0.8$
$50.9$
$\scriptscriptstyle\pm 0.5$
$62.3$
$\scriptscriptstyle\pm 0.4$
$14.0$
$\scriptscriptstyle\pm 0.3$
$31.5$
$\scriptscriptstyle\pm 0.2$
$42.9$
$\scriptscriptstyle\pm 0.2$
KIP [86, 87] KRR
$90.1$
$\scriptscriptstyle\pm 0.1$
$97.5$
$\scriptscriptstyle\pm 0.0$
$98.3$
$\scriptscriptstyle\pm 0.1$
$73.5$
$\scriptscriptstyle\pm 0.5$
$86.8$
$\scriptscriptstyle\pm 0.1$
$88.0$
$\scriptscriptstyle\pm 0.1$
$\mathbf{57.3}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$75.0$
$\scriptscriptstyle\pm 0.1$
$\mathbf{85.0}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{49.9}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
$62.7$
$\scriptscriptstyle\pm 0.3$
$68.6$
$\scriptscriptstyle\pm 0.2$
$15.7$
$\scriptscriptstyle\pm 0.2$
$28.3$
$\scriptscriptstyle\pm 0.1$
FRePo [132] KRR $\mathbf{93.0}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{98.6}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{99.2}$ $\scriptscriptstyle\mathbf{\pm 0.0}$ $\mathbf{75.6}$ $\scriptscriptstyle\mathbf{\pm 0.3}$
$86.2$
$\scriptscriptstyle\pm 0.2$
$\mathbf{89.6}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ - - -
$46.8$
$\scriptscriptstyle\pm 0.7$
$65.5$
$\scriptscriptstyle\pm 0.4$
$71.7$
$\scriptscriptstyle\pm 0.2$
$\mathbf{28.7}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{42.5}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
$44.3$
$\scriptscriptstyle\pm 0.2$
$\mathbf{15.4}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{25.4}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ -
RFAD [74] KRR $\mathbf{94.4}$ $\scriptscriptstyle\mathbf{\pm 1.5}$ $\mathbf{98.5}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$98.8$
$\scriptscriptstyle\pm 0.1$
$\mathbf{78.6}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{87.0}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{88.8}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{52.2}$ $\scriptscriptstyle\mathbf{\pm 2.2}$
$74.9$
$\scriptscriptstyle\pm 0.4$
$80.9$
$\scriptscriptstyle\pm 0.3$
$\mathbf{53.6}$ $\scriptscriptstyle\mathbf{\pm 1.2}$
$66.3$
$\scriptscriptstyle\pm 0.5$
$71.1$
$\scriptscriptstyle\pm 0.4$
$\mathbf{26.3}$ $\scriptscriptstyle\mathbf{\pm 1.1}$
$33.0$
$\scriptscriptstyle\pm 0.3$
IDC [54] GM - - - - - -
$68.1$
$\scriptscriptstyle\pm 0.1$
$87.3$
$\scriptscriptstyle\pm 0.2$
$90.2$
$\scriptscriptstyle\pm 0.1$
$50.0$
$\scriptscriptstyle\pm 0.4$
$67.5$
$\scriptscriptstyle\pm 0.5$
$74.5$
$\scriptscriptstyle\pm 0.1$
$44.8$
$\scriptscriptstyle\pm 0.2$
RTP [26] BPTT $\mathbf{98.7}$ $\scriptscriptstyle\mathbf{\pm 0.7}$ $\mathbf{99.3}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{99.4}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{88.5}$ $\scriptscriptstyle\mathbf{\pm 0.1}$ $\mathbf{90.0}$ $\scriptscriptstyle\mathbf{\pm 0.7}$ $\mathbf{91.2}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{87.3}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$89.1$
$\scriptscriptstyle\pm 0.2$
$89.5$
$\scriptscriptstyle\pm 0.2$
$\mathbf{66.4}$ $\scriptscriptstyle\mathbf{\pm 0.4}$
$71.2$
$\scriptscriptstyle\pm 0.4$
$73.6$
$\scriptscriptstyle\pm 0.4$
$34.0$
$\scriptscriptstyle\pm 0.4$
$42.9$
$\scriptscriptstyle\pm 0.7$
$16.0$
$\scriptscriptstyle\pm 0.7$
HaBa [71] TM - - - - - -
$69.8$
$\scriptscriptstyle\pm 1.3$
$83.2$
$\scriptscriptstyle\pm 0.4$
$88.3$
$\scriptscriptstyle\pm 0.1$
$48.3$
$\scriptscriptstyle\pm 0.8$
$69.9$
$\scriptscriptstyle\pm 0.4$
$74.0$
$\scriptscriptstyle\pm 0.2$
$33.4$
$\scriptscriptstyle\pm 0.4$
$40.2$
$\scriptscriptstyle\pm 0.2$
$\mathbf{47.0}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ - - -
KFS [62] DM - - - - - -
$82.9$
$\scriptscriptstyle\pm 0.4$
$\mathbf{91.4}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{92.2}$ $\scriptscriptstyle\mathbf{\pm 0.1}$
$59.8$
$\scriptscriptstyle\pm 0.5$
$\mathbf{72.0}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{75.0}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{40.0}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{50.6}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ - $\mathbf{22.7}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{27.8}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ -
Whole -
$99.6$
$\scriptscriptstyle\pm 0.0$
$93.5$
$\scriptscriptstyle\pm 0.1$
$95.4$
$\scriptscriptstyle\pm 0.1$
$84.8$
$\scriptscriptstyle\pm 0.1$
$56.2$
$\scriptscriptstyle\pm 0.3$
$37.6$
$\scriptscriptstyle\pm 0.4$
Table 2: Performance comparison of different dataset distillation methods on ImageNet-1K.
Methods IPC=11 IPC = 22 IPC = $10$ IPC = $50$
Random
$0.5$
$\scriptscriptstyle\pm 0.1$
$0.9$
$\scriptscriptstyle\pm 0.1$
$3.6$
$\scriptscriptstyle\pm 0.1$
$15.3$
$\scriptscriptstyle\pm 2.3$
DM [128]
$1.5$
$\scriptscriptstyle\pm 0.1$
$1.7$
$\scriptscriptstyle\pm 0.1$
FRePo [132]
$7.5$
$\scriptscriptstyle\pm 0.3$
$9.7$
$\scriptscriptstyle\pm 0.2$
TESLA [25] $\mathbf{7.7}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{10.5}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{17.8}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{27.9}$ $\scriptscriptstyle\mathbf{\pm 1.2}$
Whole
$33.8$
$\scriptscriptstyle\pm 0.3$
To demonstrate the effectiveness of dataset distillation, we collect and summarize the classification performance of some characteristic dataset distillation approaches on the following image datasets: MNIST [61], FashionMNIST [118] , SVHN [85], CIFAR-10/100 [56], Tiny ImageNet [60], and ImageNet-1K [95]. The details of these datasets are presented as follows. Recently, Cui et al. [24] publish a benchmark in terms of dataset distillation, however, it only contains five DD methods, and we provide a comprehensive comparison of over $15$ existing dataset distillation methods in this survey.
MNIST is a black-and-white dataset that consists of $60,000$ training images and $10,000$ test images from $10$ different classes; the size of each example image is $28\times 28$ pixels. SVHN is a colorful dataset and consists of $73,257$ digits and $26,032$ digits for training and testing, respectively, and the example images in SVHN are $32\times 32$ RGB images. CIFAR-10/100 are composed of $50,000$ training images and $10,000$ test images from $10$ and $100$ different classes, respectively. The RGB images in CIFAR-10/100 are comprised of $32\times 32$ pixels. Tiny ImageNet consists of $100,000$ and $10,000$ $64\times 64$ RGB images from $200$ different classes for training and testing, respectively. ImageNet-1K is a large image dataset that consists of over $1,000,000$ high-resolution RGB images from $1,000$ different classes.
An important factor affecting the test accuracy w.r.t. the distilled dataset is the distillation budget, which constrains the size of the distilled dataset by the notion of images allocated per class (IPC). Usually, the distilled dataset is set to have the same number of classes as the target dataset. Therefore, for a target dataset with $100$ classes, setting the distillation budget to $\text{IPC}=10$ suggests that there are a total of $10\times 100=1,000$ images in the distilled dataset.
Table 3: Cross-architecture performance w.r.t. ResNet-18 (RN-18) and ResNet-152 (RN-152) on CIFAR-10. The results of DC, DSA, MTT, DM, and KIP are derived from Cui et al. [24]. The results of FTD and CAFE are collected from Du et al. [29]; and the result of TESLA is derived from Cui et al. [25].
Methods IPC=11 IPC=$10$ IPC=$50$
ConvNet RN-18 RN-152 ConvNet RN-18 RN-152 ConvNet RN-18 RN-152
DC [129]
$28.3$
$\scriptscriptstyle\pm 0.5$
$25.6$
$\scriptscriptstyle\pm 0.6$
$\mathbf{15.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$
$44.9$
$\scriptscriptstyle\pm 0.5$
$42.1$
$\scriptscriptstyle\pm 0.6$
$16.1$
$\scriptscriptstyle\pm 1.0$
$53.9$
$\scriptscriptstyle\pm 0.5$
$45.9$
$\scriptscriptstyle\pm 1.4$
$19.7$
$\scriptscriptstyle\pm 1.2$
DSA [126]
$28.8$
$\scriptscriptstyle\pm 0.5$
$25.6$
$\scriptscriptstyle\pm 0.6$
$15.1$
$\scriptscriptstyle\pm 0.7$
$52.1$
$\scriptscriptstyle\pm 0.5$
$42.1$
$\scriptscriptstyle\pm 0.6$
$16.1$
$\scriptscriptstyle\pm 1.0$
$60.6$
$\scriptscriptstyle\pm 0.5$
$49.5$
$\scriptscriptstyle\pm 0.7$
$20.0$
$\scriptscriptstyle\pm 1.2$
MTT [13]
$46.3$
$\scriptscriptstyle\pm 0.8$
$\mathbf{34.2}$ $\scriptscriptstyle\mathbf{\pm 1.4}$
$13.4$
$\scriptscriptstyle\pm 0.9$
$65.3$
$\scriptscriptstyle\pm 0.7$
$38.8$
$\scriptscriptstyle\pm 0.7$
$15.9$
$\scriptscriptstyle\pm 0.2$
$71.6$
$\scriptscriptstyle\pm 0.2$
$60.0$
$\scriptscriptstyle\pm 0.7$
$20.9$
$\scriptscriptstyle\pm 1.6$
FTD [29] - - - - - - $\mathbf{73.8}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{65.7}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ -
TESLA [25] - - - $\mathbf{66.4}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ $\mathbf{48.9}$ $\scriptscriptstyle\mathbf{\pm 2.2}$ - - - -
DM [128]
$26.0$
$\scriptscriptstyle\pm 0.8$
$20.6$
$\scriptscriptstyle\pm 0.5$
$14.1$
$\scriptscriptstyle\pm 0.6$
$48.9$
$\scriptscriptstyle\pm 0.6$
$38.2$
$\scriptscriptstyle\pm 1.1$
$15.6$
$\scriptscriptstyle\pm 1.5$
$63.0$
$\scriptscriptstyle\pm 0.4$
$52.8$
$\scriptscriptstyle\pm 0.4$
$\mathbf{21.7}$ $\scriptscriptstyle\mathbf{\pm 1.3}$
CAFE [113] - - - - - -
$55.5$
$\scriptscriptstyle\pm 0.4$
$25.3$
$\scriptscriptstyle\pm 0.9$
KIP [87] $\mathbf{49.9}$ $\scriptscriptstyle\mathbf{\pm 0.2}$
$27.6$
$\scriptscriptstyle\pm 1.1$
$14.2$
$\scriptscriptstyle\pm 0.8$
$62.7$
$\scriptscriptstyle\pm 0.3$
$45.2$
$\scriptscriptstyle\pm 1.4$
$\mathbf{16.6}$ $\scriptscriptstyle\mathbf{\pm 1.4}$
$68.6$
$\scriptscriptstyle\pm 0.2$
$60.0$
$\scriptscriptstyle\pm 0.7$
$20.9$
$\scriptscriptstyle\pm 1.6$
Table 4: Cross-architecture performance w.r.t. ResNet-10 (RN-10) and DenseNet-121 (DN-121) on CIFAR-10. The results are derived from Lee et al. [62]
Methods IPC=11 IPC=$10$ IPC=$50$
ConvNet RN-10 DN-121 ConvNet RN-10 RN-121 ConvNet RN-10 RN-121
DSA [126]
$28.8$
$\scriptscriptstyle\pm 0.7$
$25.1$
$\scriptscriptstyle\pm 0.8$
$25.9$
$\scriptscriptstyle\pm 1.8$
$52.1$
$\scriptscriptstyle\pm 0.5$
$31.4$
$\scriptscriptstyle\pm 0.9$
$32.9$
$\scriptscriptstyle\pm 1.0$
$60.6$
$\scriptscriptstyle\pm 0.5$
$49.0$
$\scriptscriptstyle\pm 0.7$
$53.4$
$\scriptscriptstyle\pm 0.8$
DM [128]
$26.0$
$\scriptscriptstyle\pm 0.8$
$13.7$
$\scriptscriptstyle\pm 1.6$
$12.9$
$\scriptscriptstyle\pm 1.8$
$48.9$
$\scriptscriptstyle\pm 0.6$
$31.7$
$\scriptscriptstyle\pm 1.1$
$32.2$
$\scriptscriptstyle\pm 0.8$
$63.0$
$\scriptscriptstyle\pm 0.4$
$49.1$
$\scriptscriptstyle\pm 0.7$
$53.7$
$\scriptscriptstyle\pm 0.7$
IDC [54]
$50.0$
$\scriptscriptstyle\pm 0.4$
$41.9$
$\scriptscriptstyle\pm 0.6$
$39.8$
$\scriptscriptstyle\pm 1.2$
$67.5$
$\scriptscriptstyle\pm 0.5$
$63.5$
$\scriptscriptstyle\pm 0.1$
$61.6$
$\scriptscriptstyle\pm 0.6$
$74.5$
$\scriptscriptstyle\pm 0.1$
$72.4$
$\scriptscriptstyle\pm 0.5$
$71.8$
$\scriptscriptstyle\pm 0.6$
KFS [62] $\mathbf{59.8}$ $\scriptscriptstyle\mathbf{\pm 0.5}$ $\mathbf{47.0}$ $\scriptscriptstyle\mathbf{\pm 0.8}$ $\mathbf{49.5}$ $\scriptscriptstyle\mathbf{\pm 1.3}$ $\mathbf{72.0}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{70.3}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{71.4}$ $\scriptscriptstyle\mathbf{\pm 0.4}$ $\mathbf{75.0}$ $\scriptscriptstyle\mathbf{\pm 0.2}$ $\mathbf{75.1}$ $\scriptscriptstyle\mathbf{\pm 0.3}$ $\mathbf{76.3}$ $\scriptscriptstyle\mathbf{\pm 0.4}$
6.1 Standard Benchmark
For a comprehensive comparison, we also present the test accuracy w.r.t. a random select search, coreset approach, and the original target dataset. For the coreset approach, the Herding algorithm [94] is employed in this survey due to its superior performance. Notably, for most of parametric DD methods, ConvNet [33] is the default architecture to distill the synthetic data; and all the test accuracy of distilled data are obtained by training the same architecture of ConvNet for fair comparison.
The empirical comparison results associated with MNIST, FashionMNIST, SVHN, CIFAR-10/100, and Tiny ImageNet are presented in Table 1. In spite of severe scale issues and few implementations on ImageNet-1K, we present the DD results of ImageNet-1K in Table 2 for completion. As shown in these tables, many methods are not tested on some datasets due to the lack of meaning for easy datasets or scalability problems for large datasets. We preserve these blanks in tables for a more fair and comprehensive comparison. To make the comparison clear, we use the orange number for the best of two nonfactorized DD methods and blue number for the best factorized DD methods in each category in Table 1. In addition, we note the factorized DD methods with . Based on the performance comparison in the tables, we observe the followings:
  • Dataset distillation can be realized on many datasets with various sizes.
  • Dataset distillation methods outperforms random pick and coreset selection by large margins.
  • KRR and trajectory matching approaches possess advanced DD performance w.r.t. three larger datasets of CIFAR-10/100 and Tiny ImageNet among nonfactorized DD methods.
  • Factorized dataset distillation by optimizing the latent codes and decoders can largely improve the test accuracy of distilled data.
  • KFS [62] has the best overall performance among different datasets of CIFAR-10/100 and Tiny ImageNet.
6.2 Cross-Architecture Transferability
The standard benchmark in Table 1 only shows the performance of DD methods in terms of the specific ConvNet, while the distilled dataset is usually used to train other unseen network architectures. Therefore, measuring the DD performance on different architecture is also important for more comprehensive evaluation. However, there is not a standard set of architectures to evaluate cross-architecture transferability of DD algorithms, and the choice of architectures are highly different in the DD literature.
In this survey, we collect the DD results on four popular architectures of ResNet-10/18/152 [41] and DenseNet-121 [46] to evaluate the cross-architecture transferability, as shown in Tables 3 and 4. By investigating these tables, we can obtain several observations as follows:
  • There is a significant performance drop for nonfactorized DD methods when the distilled data are used to train unseen architectures (Table 3).
  • The distilled data are not always possess the best performance across different architectures, which underlines the importance of evaluation on multiple architectures (Table 3).
  • The factorized DD methods (IDC and KFS) possess better cross-architecture transferasbility i.e., suffer a smaller accuracy drop when encounter unseen architectures (Table 4).
7 Application
Due to this superior performance in compressing massive datasets, dataset distillation has been widely employed in many application domains that limit training efficiency and storage, including continual learning and neural architecture search. Furthermore, due to the correspondence between examples and gradients, dataset distillation can also benefit privacy preservation, federated learning, and adversarial robustness. In this section, we briefly review these applications w.r.t dataset distillation.
7.1 Continual Learning
During to the training process, when there is a shift in the training data distribution, the model will suddenly lose its ability to predict the previous data distribution. This phenomenon is referred to as catastrophic forgetting and is common in deep learning. To overcome this problem, continual learning has been developed to incrementally learn new tasks while preserving performance on old tasks [94, 12]. A common method used in continual learning is the replay-based strategy, which allows a limited memory to store a few training examples for rehearsal in the following training. Therefore, the key to the replay-based strategy is to select highly informative training examples to store. Benefiting from extracting the essence of datasets, the dataset distillation technique has been employed to compress data for memory with limited storage [129, 11]. Because the incoming data have a changing distribution, the frequency is high for updating the elements in memory, which leads to strict requirements for the efficiency of dataset distillation algorithms. To conveniently embed dataset distillation into the replay-based strategy, Wiewel and Yang [116] and Sangermano et al. [99] decomposed the process of synthetic data generation by linear or nonlinear combination, and thus fewer parameters were optimized during dataset distillation. In addition to enhancing replay-based methods, dataset distillation can also learn a sequence of stable datasets, and the network trained on these stable datasets will not suffer from catastrophic forgetting [79].
7.2 Neural Architecture Search
For a given dataset, the technique of neural architecture search (NAS) aims to find an optimal architecture from thousands of network candidates for better generalization. The NAS process usually includes training the network candidates on a small proxy of the original dataset to save training time, and the generalization ranking can be estimated according to these trained network candidates. Therefore, it is important to design the proxy dataset such that models trained on it can reflect the model’s true performance in terms of the original data, but the size of the proxy dataset requires control for the sake of efficiency. To construct proxy datasets, conventional methods have been developed, including random selection or greedy search, without altering the original data [9, 67]. Such et al. [105] first proposed optimizing a highly informative dataset as the proxy for network candidate selection. More subsequent works have considered NAS as an auxiliary task for testing the proposed dataset distillation algorithms [129, 126, 128, 29, 24]. Through simulation on the synthetic dataset, a fairly accurate generalization ranking can be collected for selecting the optimal architecture while considerably reducing the training time.
7.3 Privacy Protection
As overparameterized neural networks easily memorize all the training data, there exists the risk for privacy leakage by inferring the well-trained networks [101, 73, 98, 44]. With a synthetic dataset, the network avoids explicitly training on the original dataset and consequently helps protect data privacy [132]. In addition, Chen et al. [18] employed gradient matching to distil private datasets by adding differential privacy noise to the gradients of synthetic data. Theoretically, Dong et al. [28] and Zheng and Li [130] built the connection between dataset distillation and differential privacy, and proved the superiority of dataset distillation in privacy preservation over conventional private data generation methods. However, Carlini et al. [10] argued that there exists flaws in both experimental and theoretical evidences in Dong et al. [28]: (1) membership inference baseline should include all training examples, while only $1\%$ of training points are used to match the synthetic sample size, which thus contributes to the high attack accuracy of baseline; and (2) the privacy analysis is based on a assumption stating that the output of learning algorithms follows a exponential distribution in terms of loss; and the assumption has already implied differential privacy.
7.4 Federated Learning
Federated learning has received increasing attention during training neural networks in the past few years due to its advantages in distributed training and private data protection [120]. The federated learning framework consists of multiple clients and one central server, and each client possesses exclusive data for training the corresponding local model [80]. In one round of federated learning, clients transmit the induced gradients or model parameters to the server after training with their exclusive data. Then the central server aggregates the received gradients or parameters to update the model parameters and broadcast new parameters to clients for the next round. In the federated learning scenario, the data distributed in different clients are often non-i.i.d data, which causes a biased minimum w.r.t. the local model and significantly hinders the convergence speed. Consequently, the data heterogeneity remarkably imposes a communication burden between the server and clients. Goetz and Tewari [34] proposed distilling a small set of synthetic data from original exclusive data by matching gradients, and then the synthetic data instead of a large number of gradients were transmitted to the server for model updating. Other distillation approaches such as BPTT [45, 131], KRR [103], and distribution matching [119], were also employed to compress the exclusive data to alleviate the communication cost at each transition round. However, Liu et al. [70] discovered that synthetic data generated by dataset distillation are still heterogeneous. To address the problem, these researchers proposed two strategies of dynamic weight assignment and meta knowledge sharing during the distillation process, which significantly accelerate the convergence speed of federated learning. Apart from compressing the local data, Pi et al. [90] also distilled data via trajectory matching in the server, which allows the synthetic data to possess global information. Then the distilled data can be used to fine-tune the server model for convergence speed up.
7.5 Adversarial Robustness
Deep learning networks with standard training are fragile with respect to adversarial attacks, and adding imperceptible perturbations to the input can completely fool the network [37]. Adversarial training has been widely used to address this problem by continuously feeding adversarial examples during the training process [78]. However, the generation of adversarial examples requires multistep gradient ascent, which considerably undermines training efficiency. Recently, Tsilivis et al. [111] and Wu et al. [117] employed dataset distillation to extract the information of adversarial examples and generate robust datasets. Then standard network training on the distilled robust dataset is sufficient to achieve satisfactory robustness to adversarial perturbations, which substantially saves computing resources compared to expensive adversarial training.
7.6 Other Applications
Apart from the aforementioned applications, we also summarize other applications in terms of dataset distillation as follows.
Because of their small size, synthetic datasets have been applied to explainability algorithms [74]. In particular, due to the small size of synthetic data, it is easy to measure how the synthetic examples influence the prediction of test examples. If the test and training images both rely on the same synthetic images, then the training image will greatly influence the prediction of the test image. In other words, the synthetic set becomes a bridge to connect the training and testing examples.
Using the characteristic of capturing the essence of datasets, [14, 21] utilized dataset distillation for visual design. Cazenavette et al. [14] generated the representative textures by randomly cropping the synthetic images during the distillation process. In addition to extracting texture, Chen et al. [21] imposed the synthetic images to model the outfit compatibility through dataset distillation.
In addition to the continuous image domain, dataset distillation has also been extended to discrete data such as graphs [50, 49, 69] and recommender systems [96]. Jin et al. [50] first formulated this dataset distillation w.r.t. graph data, and successfully distilled a small, condensed graph via gradient matching. Due to the computational inefficiency in distilling pairwise relations for graph nodes, Jin et al. [49] turned to learning a probabilistic graph model, which allows a differentiable optimization w.r.t. the discrete graph. These authors also employed one-step strategy for further distillation speedup. Moreover, Liu et al. [71] accelerated the graph distillation through distribution matching from the perspective of receptive fields. For the recommender system, Sachdeva et al. [96] distilled a massive dataset to a continuous prior for tackling the discrete data problem.
Due to their small size and abstract visual information, distilled data can also be applied in medicine, especially in medical image sharing [66]. Empirical studies on gastric X-ray images have shown the advantages of DD in medical image transition and the anonymization of patient information [64, 65]. Through dataset distillation, hospitals can share their valuable medical data at lower cost and risk to collaboratively build powerful computer-aided diagnosis systems.
For backdoor attack, Liu et al. [72] considered to inject backdoor triggers into the small distilled dataset. Because the synthetic data has a small size, directly trigger insertion is less effective and also perceptible. Therefore, these coworkers turned to insert triggers into the target dataset during the dataset distillation procedure. In addition, they propose to iteratively optimize the triggers during the distillation process to preserve the triggers’ information for better backdoor attack performance.
8 Challenges and Directions
Due to its superiority in compressing datasets and training speedup, dataset distillation has promising prospects in a wide range of areas. In the following, we discuss existing challenges in terms of dataset distillation. Furthermore, we investigate plausible directions and provide insights into dataset distillation to promote future studies.
8.1 Challenges
For dataset distillation, there are two main ingredients: (1) measure the difference between the knowledge of synthetic data and real (original) data; and (2) update the synthetic data to narrow the knowledge difference. Based on these ingredients, we discuss the challenges of dataset distillation from the perspectives of (1) the definition of the difference in the knowledge between two types of datasets, (2) the form of synthetic datasets, (3) the evaluation of synthetic datasets, and (4) the theory behind dataset distillation.
Dataset distillation methods aim to transfer the knowledge of a target dataset into a small synthetic dataset to achieve a comparable generalization with respect to the original data distribution. Although the knowledge of datasets is an abstract notion, many DD approaches indirectly measure the difference in the knowledge base via various proxy losses. However, the existing definition w.r.t. the knowledge difference is inconsistent with the original comparable generalization. For example, the meta-loss $\mathcal{R}_{\mathcal{T}}(\texttt{alg}(\mathcal{S}))$ in the meta-learning framework measure the knowledge difference in the error in terms of a specific target dataset. Through this, one can only obtain a similar training error other than the test error w.r.t. the target dataset. As for the data matching framework, it intuitively makes knowledge concrete with a series of parameters or features, which are matched for knowledge transmission. Therefore, a more consistent definition of knowledge difference should be investigated to improve the dataset distillation performance.
Although most dataset distillation methods directly update the synthetic images, optimizing factorized codes and decoders that cooperatively generate synthetic data outperform the direct optimization by a large margin, as discussed in Section 6.1. From this, the predefined form of the code and decoder has a remarkable influence on the distillation performance. Many Factorization methods on synthetic data are based on intuition and lack rigorous analysis, and only some vague reasons are proposed, such as decreasing redundant information across different classes and learning more compact representations. Hence, it is necessary to explore the influence of the predefined factorization on the dataset distillation for learning a more informative distilled dataset.
In practical dataset distillation, the optimization of synthetic data is closely related to the specific network architecture (or kernels in KRR approach), and there is naturally a performance drop when the learned synthetic dataset is applied to other unseen architectures; please see Section 6.2. Therefore, it is one-sided to compare different DD methods only on one or a few architectures. To compare different DD methods in a more comprehensive scenario, it is essential to set a reasonable baseline consisting of multiple handpicked network architectures with different complexity.
Although various DD approaches have been emerging, there have been few investigations that discuss the theory behind dataset distillation. Nevertheless, developing the theory is extremely necessary and will considerably promote the dataset distillation by directly improving the performance and also suggesting the correct direction of development. For example, the theory can help derive a better definition of the dataset knowledge and consequently increase the performance. In addition, the theoretical correlation between the distillation budget (the size of synthetic datasets) and performance can provide an upper bound on the test error, which can offer a holistic understanding and prevent researchers from blindly improving the DD performance. Hence, a solid theory is indispensable to take the development of DD to the next stage.
8.2 Future Directions
Although the existing DD methods have shown impressive results in compressing massive datasets, there are still many areas that are worth exploring to promote dataset distillation. In this section, we will discuss some promising directions that shed light on future studies.
Fine-tune with synthetic data. With the rapid development of foundational models [7], the mainstream deep learning pipeline gradually converts from scratch learning to pretrained learning or transfer learning. The foundational models of BERT [27], GPT [92, 93, 8], etc., are trained with billions of data points to learn an appropriate latent representation. With only fine-tuning on specific datasets, the foundational models can achieve extraordinary performance on various downstream tasks and outperform models learned from scratch by a large margin. To this end, conducting dataset distillation in the fine-tuning regime is a promising direction to cooperate with foundational models. With synthetic data, the fine-tuning process will be faster while the downstream performance is preserved. Despite possessing a similar goal, distilling datasets based on pretrained models might be more challenging than the initialized models because the pretrained models are harder to collect for repeated distillation to enhance synthetic datasets’ robustness w.r.t. model initialization. Apart from condensing datasets in the fine-tuning stage, it is also an interesting topic to distil datasets in the vanilla training stage so that models trained on synthetic datasets can be easier or faster to fine-tune by specific datasets.
Initialization augmentation. Benefiting from excellent efficiency and compatibility, data augmentation has become an essential strategy for enhancing the model’s generalization w.r.t. data distribution [102, 15, 88]. From basic cropping and flipping [121] to the mixup method [123], data augmentation efficiently constructs variants of training images with prior rules to improve the model’s robustness to patterns. In dataset distillation, synthetic datasets are also sensitive to the parameter initialization and network architecture adopted in the distillation process [114, 129]. Existing works alleviate this problem by trivial repeated distillation w.r.t. different initializations and then taking an expectation [132]. However, the random selection of initialization is not an elaborate strategy and has limited improvement. Zhang et al. [124] showed that employing early-stage models trained with a few epochs as the initialization can achieve better distillation performance, and they further proposed weight perturbation methods to efficiently generate early-stage models for repeated distillation. Therefore, it is important for many dataset distillation algorithms to investigate how to design the initialization augmentation so that synthetic datasets achieve better generalization to different initializations and architectures.
Preprocessing of target datasets. To better extract knowledge from training data, preprocessing, such as normalization, is a common approach that reshapes the structure of data by well-designed mappings [32]. In existing dataset distillation methods, only a few works discuss the influence of preprocessing on final distillation performance [86]. Nguyen et al. [86, 87] and Cazenavette et al. [13] preprocessed target datasets with zero-phase compenent analysis (ZCA) whitening [53] before distillation and achieved satisfying empirical results. Through elaborate preprocessing, more explicit information can emerge, and the target datasets might be easier to distill. In addition, this preprocessing is compatible with existing dataset distillation algorithms and thus can be deployed without extra effort. For this reason, it is useful to investigate the preprocessing of target datasets for distillation acceleration and also performance improvement.
Acknowledge
We sincerely thank Awesome-Dataset-Distillation for its comprehensive and timely DD publication summary. We also thank Can Chen and Guang Li for their kind feedback and suggestions.